A firm has a market capitalization (market value of equity) of $25 Billion and net debt of $15 Billion. Calculate the weight of debt in the firm's weighted average cost of capital (WACC) calculation. [Note: Enter your answer as a percentage rounded to two decimal places.]
Answer #1

Total firm value= $25 + $15= $40

Weight of debt in weighted average cost of capital:

= $15/ $40

= 0.3750*100

= 37.50%.

In case of any query, kindly comment on the solution.
Answer #2

ANSWER :




C = E + D = 25 + 15 = 40 billion dollars.


Weight of debt in the capital = D/C = 15 / 40 = 0.375 = 37.50% (ANSWER).

answered by: Tulsiram Garg